If you know the names of the remaining six students in the spelling bee, what is the probability of randomly selecting an order and getting the order that is used in the spelling bee?
Answer #1

Total students = 6

Number of ways to order them = 6! = 720

Probability of getting selected order = 1/6! = 1/720 = 0.001388